OpenStudy (anonymous):

PLEASE HELP, IM LOST ON THIS QUESTION Given the equation 2Square root of x minus 5 = 2, solve for x and identify if it is an extraneous solution.

OpenStudy (anonymous):

@@peachpi @imqwerty @arindameducationusc @godgirl122 @heretohelpalways please help I'm confused

OpenStudy (anonymous):

i did already but i got 2x-5=4

OpenStudy (anonymous):

or would it be 4x-5=4?

imqwerty (imqwerty):

is this the question - \[\sqrt[2]{x-5}=2 \]??

OpenStudy (anonymous):

yes

imqwerty (imqwerty):

ok so when u square both the sides \[(\sqrt[2]{x+5})^2 =2^2\] what will be the RHS after squaring? :)

OpenStudy (anonymous):

sorry , whats the RHS again?

imqwerty (imqwerty):

RHS=right hand side :)

OpenStudy (anonymous):

ok thanks it would be

imqwerty (imqwerty):

right hand side of the equation

OpenStudy (anonymous):

4

imqwerty (imqwerty):

yes and what wuld be the Left hand side?

OpenStudy (anonymous):

x+5

imqwerty (imqwerty):

so we get x+5 =4

OpenStudy (anonymous):

yes, now we solve for the variable and get -1 right?

OpenStudy (anonymous):

@imqwerty

OpenStudy (anonymous):

please help i don't know what to do after this step

imqwerty (imqwerty):

yes u get x=-1 :)

OpenStudy (anonymous):

@imqwerty wait don't go please help

imqwerty (imqwerty):

ok

imqwerty (imqwerty):

u need help in??

OpenStudy (anonymous):

well my answer choices are these : x= 6, solution is not extraneous x = 6, solution is extraneous x = 11, solution is not extraneous x = 11, solution is extraneous

OpenStudy (anonymous):

and none of those are -1 so i don't know what to do next

OpenStudy (anonymous):

@imqwerty @peachpi

imqwerty (imqwerty):

wait

imqwerty (imqwerty):

we got x-5 = 1 add 5 to both sides u get x=6

OpenStudy (anonymous):

ok thank you so much

OpenStudy (anonymous):

how do i know if it is extraneous or not

imqwerty (imqwerty):

and the solution is not extraneous cause when u put x=6 is a true solution :)

OpenStudy (anonymous):

ok so if its a true solution its always not extraneous?

imqwerty (imqwerty):

yes :)

OpenStudy (anonymous):

ok, thank you for all of your help! it was much appreciated

imqwerty (imqwerty):

no prblem :)